# GiBy: A Giant-Step Baby-Step Classifier For Anomaly Detection In Industrial Control Systems

### v1

@@ Abstract
The continuous monitoring of the interactions between cyber-physical components of any industrial control system (ICS) is required to secure automation of the system controls, and to guarantee plant processes are fail-safe and remain in an acceptably safe state. Safety is achieved by managing actuation (where electric signals are used to trigger physical movement), dependent on corresponding sensor readings; used as ground truth in decision making. Timely detection of anomalies (attacks, faults and unascertained states) in ICSs is crucial for the safe running of a plant, the safety of its personnel, and for the safe provision of any services provided. We propose an anomaly detection method that involves accurate linearization of the non-linear forms arising from sensor-actuator(s) relationships, primarily because solving linear models is easier and well understood. Further, the time complexity of the anomaly detection scenario/problem at hand is lowered using dimensionality reduction of the actuator(s) in relationship with a sensor. We accomplish this by using a well-known water treatment testbed as a use case. Our experiments show millisecond time response to detect anomalies and provide explainability; that are not simultaneously achieved by other state of the art AI/ML models with eXplainable AI (XAI) used for the same purpose. Further, we pin-point the sensor(s) and its actuation state for which anomaly was detected.
Index Terms:
Anomaly Detection, Cyber-Attacks, Safety, Security, Cyber-Physical System, Industrial Control System.I Introduction
@@
Physical/mechanical automation components are more recently becoming digitized and analogue systems are also being developed with digital interfaces to enable integration with digital systems. Industries such as manufacturing (including assembly, quality check and distribution) and critical infrastructure (CI) - including the water sector (water and wastewater treatment, and distribution), energy sector (generation, transmission and distribution), and transportation sector, play an important role in global society. The provision of high-quality CI services in particular should be noted as significant for public health, well-being and safety. The integration and embedding of computational processing units expose the ICS used in industries and CI to an expanded attack surface, making it vulnerable to cyber-attacks [1]. For example, malware was found in the manufacturing sector focused on a steel plant [2]. This attack led to an unregulated furnace and the resulting dangerous infrastructure could not be shut down as normal, leading to physical damage to the plant [3]. Recent power sector-focused malware such as FrostyGoop [4] were able to infiltrate CI and disrupt services. Disruption and damage have been recorded in the water sector due to attacks on CPS at the Maroochy Shire plant in Australia [5], multiple plants in Texas, USA [6], Mayo in Ireland [7], and a mitigated attack at the Oldsmar plant in USA [8]. The spate of real-world attacks on manufacturing and CI operational technology (OT), and their consequences reiterate the value in timely detection of these attacks and subsequent actions to bring these systems into a safe and non-compromised state.
@@
In this work, we use the Secure Water Treatment (SWaT) [9] testbed as a use case in ICS anomaly detection. We use system-specific information to improve anomaly detection, as opposed to other models that employ specialized solutions on a subset of available information [10, 11, 12], or prefer to use a generic data-driven approach [13]. Specifically, we rely on i.) the real-time data logs to retrieve state information from plant sensors and actuators, ii.) the PLC control logic sheet to determine sensor boundary conditions and corresponding actuation steps (if any), and iii.) infer sensor-actuator(s) relationships from the plant system architecture.
@@
Going over the complete actuation state space is a case of solving problems in exponential time (see Section II-B). We employ a dimensionality reduction technique to reduce the time complexity by considering only the nearest-neighbour actuators related to a given sensor. This is followed by an accurate linearization of a reduced number of actuator(s)-sensor relationships. We note that the water treatment plant (see Fig. 3), has limited nonlinear sensor-actuator(s) relationships. This is attributed to the (mostly) serial nature of numerous industrial processes, including those involving assembly lines in a factory setting and a water treatment plant. Once the sensor-actuator(s) relationships are linearized, it is trivial to determine the bounds required to detect anomalies in the ICS. The advantage of this approach is that it typically allows detection and explanation within a millisecond (see Section V). It saves valuable time with fast detection and explanation, helping with faster diagnostics, mitigation, or recovery. The novelty being that our solution is fit for use in near real-time and resource-constrained decision-making control systems (see Section VI-F). We make the following main contributions.
@@
The paper is organised as follows. Section II provides the problem context and discusses the threat model. Section III presents the detector solution. The solution is extended in Section IV. Experiments are carried out and results presented in Section V. The work is discussed in Section VI. The related work is provided in Section VII. The conclusions are drawn, and future work is proposed in Section VIII.
II Problem Context and Threat Model
II-A Sensor readings and boundary conditions
@@
Transportation components from reputable manufacturers are supplied with detailed specifications, defining the acceptable safe operational range and conditions of use. When used as part of a system, a new appropriate system-imposed limit may apply depending on the system design and other operational bottlenecks. For example, a pipe may be Original Equipment Manufacturer (OEM) safety rated for a pressure of 0-5 Bars. However, downstream constraints may only allow a pressure of 1-3 Bars. In such cases, a system-imposed limit is put in place. With respect to service provisioning boundaries, they are deduced from extensive knowledge and tests to determine the safe range for consumption. For example, a water treatment process must ensure that adding acid or base chemicals do not significantly result in the pH of water moving away from 7 (neutral).
The OEM/system imposed limits for transport components and deduced limits for service provisioning parts are fed into one or more controllers, such as a PLC (see Fig. 1). The min-to-max operational range values on the PLC assists with taking action and (possibly) alerting an operator, when the safety threshold is breached. Additionally, the PLCs may also include logic to act, even before a breach, by providing a warning buffer. This may be achieved by setting tighter boundary conditions. Traditionally, at least two layers of safety are provided, the lower boundary is the warning boundary which alerts operators to take action, the upper boundary is the min-max boundary, and this should never be passed. In some systems if the min-max boundary is passed the personnel must evacuate the site and services are put on hold until safe conditions are assured.
II-B Reducing time complexity: core idea
@@
Let there be normal and attack plant operation datasets for the sensor and actuator states logged during plant operations, under normal and attack conditions, respectively. A normal training dataset comprises of all seen (normal) state relations embedded into it. We implicitly rely on all actuator relations ) w.r.t. the sensor, but explicitly consider the actuation sequences for , for input side linearization. The excluded actuators are indirectly responsible for the change in measurement readings at sensor . However, these relations are captured at the nearest neighbour (nn-) actuators111Henceforth, it is referred to as nn-actuators. It is common to place the actuator immediately preceding the related sensor. If not, we yet consider that directly related actuator as a nearest neighbour — to improve detector resolution, when that information is available. Resolution is further enhanced by picking non-neighbour PLC imposed actuation, affecting the sensor reading; as part of the design (see Section VI-D). , and mapped to sensor readings via ) = , provided in the normal dataset. Note, only the input actuator state space is reduced. The sensor reading remains a function of all related actuators. Sensor reading is readily available from the normal dataset and its measured range, w.r.t. actuation remains unchanged. In short, we refined our search space to the sensor’s nn-actuator(s). However, the set of normal operations seen in the normal dataset might be a tiny subset of the entire normal state space. In this context, everything not seen as part of the normal plant operation dataset and derivations are treated as anomalous (similar to the concept used in one-class classification [15, 16]), and our detector flags it as a warning. The detector warnings provide an additional layer of safety along with the PLC boundaries discussed in section II-A, by generating a suitable alert when action is required.
II-C SWaT Water Treatment Testbed and datasets
@@
The SWaT water treatment (see Fig. 3) is a six-stage process ( to ) and emulates a typical treatment plant. The water for treatment is the input to and treated water is the main output of . Different (and redundant) sensors are strategically placed to make measurements across the testbed. Actuators are used for control. The sensors and actuators used in appear in Table I. Process is designed to get the raw water into the system. is focused on chemical dosing and with ultra-filtration. includes UV-dechlorination and carries out reverse osmosis (RO). uses a backwash tank for cleaning and stores treated water in another tank (that is recycled). The device names consist of two parts where the first 3 characters are shortened for devices type and the last 3 numbers refer to the process stage and number of devices. For example, the device FIT-101 refers to the flow meter in the first process where it is the first flow meter. This description of each device can be found as follows: FIT-XXX: Flow meter, LIT-XXX: Level Transmitter, AIT-XXX: Analyser, DPIT-XXX: Differential pressure indicating transmitter, PIT-XXX: Pressure meter, MV-XXX: Motorized valve, P-XXX: Pump and UV-XXX: UV Dechlorinator [9]. The dataset was collected for eleven days, of which the first seven days is under normal working conditions. The final four days included injecting/ swapping with anomalous data, involving 41 CPS attacks on the SWaT water testbed. The plant was run from an empty to a fully operational state. The attacks were conducted by altering the OT network traffic, spoofing the sensor values, and issuing bogus SCADA commands.
@@
The SWaT datasets provided the state view of the fifty-one sensors and actuators in the plant at one second resolution, as time series data. A state may include actuators in either open/close, on/off or open/transition/close state, depending on the type of the actuator used. The normal dataset consisted of 495,000 records of plant state under normal operations. The attack dataset had 449,919 records. Note that the attack dataset is a mix of normal and attack records. The attack records accounted for where less than 6% of total data [12]. If the plant was under attack during the record generation (within the dataset) then this record is tagged as Attack, and otherwise as Normal.
II-D Threat Model
@@
An attacker is assumed to have followed an attack vector [17] to gain access into plant operational technology. This adversary is assumed to be able to inject control commands and modify sensor readings and actuator status in OT. This leads to four attack types [9] — single stage single point (SSSP) where exactly one sensor/actuator data is under attack, single stage multi-point (SSMP) where multiple sensor/actuator data in exactly one process is under attack, multistage single point (MSSP) where exactly one sensor/actuator data in multiple processes are under attack, and multistage multi-point (MSMP) where multiple sensor/actuator data in multiple processes are under attack. It is assumed there are no attacks on the plant when the normal dataset is collected, which would be used as the training dataset.
III Giant-Step Baby-step Solution
III-A Design rationale and considerations
@@
The sensor and actuator values in the training dataset are viewed as ground truth data. Our core requirements [18] are (R1) near real-time detection, (R2) anomaly cause identification, and (R3) minimising false alarms. An automated response is not a core consideration because governance considerations involve human intervention and human-in-the-loop solutions [19, 20] to be qualified as compliant to safety standards i.e. a fully automation solution might not be viable. While the ability to automate is provided, the level of automation is left to the discretion of the implementer. One-class classification is used due to the typical imbalance between normal and attack dataset [21]; the latter contributing only a tiny percentage of attack data. We train using the normal dataset and employ unsupervised learning [22].
III-B Solution outline
@@
To meet the requirements in Section III-A, we use binary classification to solve a decision problem [23]. The proposed anomaly detector has two main steps — giant and baby step. The giant-step is used to determine [LB, UB] boundary for measured sensor values (available in normal dataset), in relation with every nn-actuation state. For example, consider the more detailed SWaT process stages as illustrated in Fig. 4. In this system, water flows into the tank when motorized valve (actuator) MV101 transitions from close to open. The tank level indicator (sensor) LIT101 measures the water tanks instantaneous water level. When pump (actuator) P101 is turned on, water is pumped out from the tank, and the flow transmission (sensor) FIT201 registers non-zero flow values. The two nn-actuators for sensor LIT101 are MV101 and P101. MV101 has three actuation states (close-1/transition-0/open-2) and P101 has two actuation states (on-2/off-1), in combination this allows for six possible states across the two actuators connected to the sensor LIT101. The bounds for LIT101 is computed by our giant-step solution for all actuator combinations of MV101 and P101: 11, 12, 01, 02, 21, and 22.
@@
III-C A switchboard to map actuation state, building linearized state groups, and enable explanation

### v3

@@ Abstract
The continuous monitoring of the interactions between cyber-physical components of any industrial control system (ICS) is required to secure automation of the system controls, and to guarantee plant processes are fail-safe and remain in an acceptably safe state. Safety is achieved by managing actuation (where electric signals are used to trigger physical movement), dependent on corresponding sensor readings; used as ground truth in decision making. Timely detection of anomalies (attacks, faults and unascertained states) in ICSs is crucial for the safe running of a plant, the safety of its personnel, and for the safe provision of any services provided. We propose an anomaly detection method that involves accurate linearization of the non-linear forms arising from sensor-actuator(s) relationships, primarily because solving linear models is easier and well understood. We accomplish this by using a well-known water treatment testbed as a use case. Our experiments show millisecond time response to detect anomalies, all of which are explainable and traceable; this simultaneous coupling of detection speed and explainability has not been achieved by other state of the art Artificial Intelligence (AI)/ Machine Learning (ML) models with eXplainable AI (XAI) used for the same purpose. Our methods explainability enables us to pin-point the sensor(s) and the actuation state(s) for which the anomaly was detected. The proposed algorithm showed an accuracy of 97.72% by flagging deviations within safe operation limits as non-anomalous; indicative that slower detectors with highest detection resolution is unnecessary, for systems whose safety boundaries provide leeway within safety limits.
1 Introduction
@@
††This work was carried out when the authors were employed at University of Bristol in the United Kingdom.Physical/mechanical automation components are more recently becoming digitized and analogue systems are also being developed with digital interfaces to enable integration with digital systems. Industries such as manufacturing (including assembly, quality check and distribution) and critical infrastructure (CI) - including the water sector (water and wastewater treatment, and distribution), energy sector (generation, transmission and distribution), and transportation sector, play an important role in global society. The provision of high-quality CI services in particular should be noted as significant for public health, well-being and safety. The integration and embedding of computational processing units expose the ICS used in industries and CI to an expanded attack surface, making it vulnerable to cyber-attacks [1]. For example, malware was found in the manufacturing sector focused on a steel plant [2]. This attack led to an unregulated furnace and the resulting dangerous infrastructure could not be shut down as normal, leading to physical damage to the plant [3]. Recent power sector-focused malware such as FrostyGoop [4] were able to infiltrate CI and disrupt services. Disruption and damage have been recorded in the water sector due to attacks on CPS at the Maroochy Shire plant in Australia [5], multiple plants in Texas, USA [6], Mayo in Ireland [7], and a mitigated attack at the Oldsmar plant in USA [8]. The spate of real-world attacks on manufacturing and CI operational technology (OT), and their consequences reiterate the value in timely detection of these attacks and subsequent actions to bring these systems into a safe and non-compromised state.
@@
In this work, we use the Secure Water Treatment (SWaT) [9] testbed as a use case in ICS anomaly detection. We use system-specific information to improve anomaly detection, as opposed to other models that employ specialized solutions on a subset of available information [10, 11, 12], or prefer to use a generic data-driven approach [13]. Specifically, we rely on i) the real-time data logs to retrieve state information from plant sensors and actuators, ii) the PLC control logic sheet to determine sensor boundary conditions and corresponding actuation steps (if any), and iii) infer sensor-actuator(s) relationships from the plant system architecture.
@@
Going over the complete actuation state space is a case of solving problems in exponential time (see Section 2.2). We employ a dimensionality reduction technique to reduce the time complexity by considering only the nearest-neighbour actuators related to a given sensor. This is followed by an accurate linearization of a reduced number of actuator(s)-sensor relationships. We note that the water treatment plant (see Fig. 4), has limited nonlinear sensor-actuator(s) relationships. This is attributed to the (mostly) serial nature of numerous industrial processes, including those involving assembly lines in a factory setting and a water treatment plant. Once the sensor-actuator(s) relationships are linearized, it is trivial to determine the bounds required to detect anomalies in the ICS. The advantage of this approach is that it typically allows detection and explanation within a millisecond (see Section 5). It saves valuable time with fast detection and explanation, helping with faster diagnostics, mitigation, or recovery. The use of per sensor anomaly detection boundaries makes our solution scalable to a number of mostly serial industrial processes (see Section 2.3.1) The novelty being that our solution is fit for use in near real-time and resource-constrained decision-making control systems, and at the same time provides accurate explanation and traceability. We make the following main contributions.
i.)
We present our anomaly detector in Section 3.
ii.)
An extended solution is proposed in Section 4 to detect anomalies that stand out over a longer period of time.
iii.)
We compare our solution results with AI/ML models for anomaly detection and explainable AI (see Section 5).
- iv.)
v.)
We discuss why some of the works in existing literature are not compatible for comparison with the detection accuracy of the proposed anomaly detector, and its implications for reporting experimental results in this work (see Section 6.7).
@@
The paper is organised as follows. Section 2 provides the problem context and discusses the threat model. Section 3 presents the detector solution. The solution is extended in Section 4. Experiments are carried out and results presented in Section 5. The work is discussed in Section 6. The related work is provided in Section 7. The conclusions are drawn, and future work is proposed in Section 8.
2 Problem Context and Threat Model
2.1 Sensor readings and boundary conditions
@@
Transportation components from reputable manufacturers are supplied with detailed specifications, defining the acceptable safe operational range and conditions of use. When used as part of a system, a new appropriate system-imposed limit may apply depending on the system design and other operational bottlenecks. For example, a pipe may be Original Equipment Manufacturer (OEM) safety rated for a pressure of 0-5 Bar. However, downstream constraints may only allow a pressure of 0-3 Bar. In such cases, a system-imposed limit is put in place. With respect to service provisioning boundaries, they are deduced from extensive knowledge and tests to determine the safe range for consumption. For example, a water treatment process must ensure that adding acid and/or base chemicals do not significantly result in the pH of water moving away from 7 (neutral).
The OEM/system imposed limits for transport components and deduced limits for service provisioning parts are fed into one or more controllers, such as a PLC (see Fig. 1). The min-to-max operational range values on the PLC assists with taking action and (possibly) alerting an operator, when the safety threshold is breached (see Section 6.6). Additionally, the PLCs may also include logic to act, even before a breach, by providing a warning buffer. This may be achieved by setting tighter boundary conditions. Traditionally, at least two layers of safety are provided, the lower boundary is the warning boundary which alerts operators to take action, the upper boundary is the min-max boundary, and this should never be passed. In some systems if the min-max boundary is passed the personnel must evacuate the site and services are put on hold until safe conditions are assured.
2.2 Reducing time complexity: core idea
@@
Let there be normal and attack plant operation datasets for the sensor and actuator states logged during plant operations, under normal and attack conditions, respectively. A normal training dataset comprises of all seen (normal) state relations embedded into it. We implicitly rely on all actuator relations ) w.r.t. the sensor, but explicitly consider the actuation sequences for , for input side linearization. The excluded actuators are indirectly responsible for the change in measurement readings at sensor . However, these relations are captured at the nearest neighbour (nn-) actuators111Henceforth, it is referred to as nn-actuators. It is common to place the actuator immediately preceding the related sensor. If not, we yet consider that directly related actuator as a nearest neighbour — to improve detector resolution, when that information is available. Resolution is further enhanced by picking non-neighbour PLC imposed actuation, affecting the sensor reading; as part of the design (see Section 6.4). , and mapped to sensor readings via ) = , provided in the normal dataset. Note, only the input actuator state space is reduced. The sensor reading remains a function of all related actuators (see Fig. 3). Sensor reading is readily available from the normal dataset and its measured range, w.r.t. actuation remains unchanged. In short, we refined our search space to the sensor’s nn-actuator(s). However, the set of normal operations seen in the normal dataset might be a tiny subset of the entire normal state space. In this context, everything not seen as part of the normal plant operation dataset and derivations are treated as anomalous (similar to the concept used in one-class classification [15, 16]), and our detector flags it as a warning. The detector warnings provide an additional layer of safety along with the PLC boundaries discussed in section 2.1, by generating a suitable alert when action is required.
2.3 SWaT Water Treatment Testbed and datasets
@@
The SWaT water treatment (see Fig. 4) is a six-stage process ( to ) and emulates a typical treatment plant. The water for treatment is the input to and treated water is the main output of . Different (and redundant) sensors are strategically placed to make measurements across the testbed. Actuators are used for control. The sensors and actuators used in appear in TABLE I. Process is designed to get the raw water into the system. is focused on chemical dosing and with ultra-filtration. includes UV-dechlorination and carries out reverse osmosis (RO). uses a backwash tank for cleaning and stores treated water in another tank (that is recycled). The device names consist of two parts where the first 3 characters are shortened for devices type and the last 3 numbers refer to the process stage and number of devices. For example, the device FIT-101 refers to the flow meter in the first process where it is the first flow meter. This description of each device can be found as follows: FIT-XXX: Flow meter, LIT-XXX: Level Transmitter, AIT-XXX: Analyser (3 types to measure ORP, pH snd conductivity), DPIT-XXX: Differential pressure indicating transmitter, PIT-XXX: Pressure meter, MV-XXX: Motorized valve, P-XXX: Pump and UV-XXX: UV Dechlorinator [9]. The dataset was collected for eleven days, of which the first seven days is under normal working conditions. The final four days included injecting/ swapping with anomalous data, involving 41 CPS attacks on the SWaT water testbed. The plant was run from an empty to a fully operational state. The attacks were conducted by altering the OT network traffic, spoofing the sensor values, and issuing bogus SCADA commands.
@@
The SWaT datasets provided the state view of the fifty-one sensors and actuators in the plant at one second resolution, as time series data. A state may include actuators in either open/close, on/off or open/transition/close state, depending on the type of the actuator used. The normal dataset consisted of 495,000 records of plant state under normal operations. The attack dataset had 449,919 records. Note that the attack dataset is a mix of normal and attack records. The attack records accounted for were less than 6% of total data [12]. If the plant was under attack during the record generation (within the dataset) then this record is tagged as Attack, and otherwise as Normal.
2.3.1 Sensor boundary independence in SWaT
SWaT has seven types of sensors — to measure tank level, flow rate, ORP, conductivity, pH, pressure and differential pressure. An anomaly detector may be designed to individually test the operational boundaries on the sensor within a process; without attempting to correlate its readings with other sensors in the same or different processes. Therefore, though the processes are interdependent, the individual sensor safety boundaries may be determined independently from its normal operation readings. When the sensor value is within determined boundaries, it is deemed safe w.r.t. the metric it measures, within the SWaT process. Hence, it is sufficient to look at the local causal paths to determine the sensor boundaries. When a sensor reading was spoofed to read within normal bounds but was actually out of bounds, we rely on other individual related sensors (if any) to report surpassing its boundary threshold. This notion of per sensor boundary computation makes the proposed anomaly detector scalable to large (mostly) serial systems. Any sensor noise will be considered as part of normal sensor behavior (provided they are within OEM tolerance limits). The data read out by the sensor itself may be linear or nonlinear. We are only interested to capture its normal operations bounds. ICS behavior may vary over long production cycles (spanning months or years) with maintenance windows and sensor drift. These are out of scope of this work.
2.4 Threat Model
@@
An attacker is assumed to have followed an attack vector [17] to gain access into plant OT. This adversary is assumed to be able to inject control commands and modify sensor readings and actuator status in OT. This leads to four attack types [9] — single stage single point (SSSP) where exactly one sensor/actuator data is under attack, single stage multi-point (SSMP) where multiple sensor/actuator data in exactly one process is under attack, multistage single point (MSSP) where exactly one sensor/actuator data in multiple processes are under attack, and multistage multi-point (MSMP) where multiple sensor/actuator data in multiple processes are under attack. It is assumed there are no attacks on the plant when the normal dataset is collected, which would be used as the training dataset.
3 Giant-Step Baby-step Solution
3.1 Design rationale and considerations
@@
The sensor and actuator values in the training dataset are viewed as ground truth data. Our core requirements [18] are (R1) near real-time detection, (R2) anomaly cause identification, and (R3) minimising false alarms. An automated response is not a core consideration because governance considerations involve human intervention and human-in-the-loop solutions [19, 20] to be qualified as compliant to safety standards i.e. a fully automated solution might not be viable. While the ability to automate is provided, the level of automation is left to the discretion of the implementer. One-class classification is used due to the typical imbalance between normal and attack dataset [21]; the latter contributing only a tiny percentage of attack data. We train using the normal dataset and employ unsupervised learning [22].
3.2 Solution outline
@@
To meet the requirements in Section 3.1, we use binary classification to solve a decision problem [23]. The proposed anomaly detector has two main steps — giant and baby step. The giant-step is used to determine [LB, UB] boundary for measured sensor values (available in normal dataset), in relation with every nn-actuation state. For example, consider the more detailed SWaT process stages as illustrated in Fig. 5. In this system, water flows into the tank when motorized valve (actuator) MV101 transitions from close to open. The tank level indicator (sensor) LIT101 measures the water tanks instantaneous water level. When pump (actuator) P101 is turned on, water is pumped out from the tank, and the flow transmission (sensor) FIT201 registers non-zero flow values. The two nn-actuators for sensor LIT101 are MV101 and P101. MV101 has three actuation states (close-1/transition-0/open-2) and P101 has two actuation states (on-2/off-1), in combination this allows for six possible states across the two actuators connected to the sensor LIT101. The bounds for LIT101 is computed by our giant-step solution for all actuator combinations of MV101 and P101: 11, 12, 01, 02, 21, and 22.
@@
3.3 A switchboard to map actuation state, building linearized state groups, and enable explanation